# Online Energy Storage Arbitrage under Imperfect Predictions: A Conformal Risk-Aware Approach

### v1

@@ Abstract.
This work proposes a conformal approach for energy storage arbitrage to control the downside risks arose from imperfect price forecasts. Energy storage arbitrage relies solely on predictions of future market prices, while inaccurate price predictions may lead to significant profit losses. Based on conformal decision theory, we describe a controller that dynamically adjusts decision conservativeness through prediction sets without distributional assumptions. To enable online calibration when online profit loss feedback is unobservable, we establish that a temporal difference error serves as a measurable proxy. Building on this insight, we develop two online calibration strategies: prediction error-based adaptation targeting forecast accuracy, and value error-based calibration focusing on decision quality. Analysis of the conformal controller proves bounded long-term risk with convergence guarantees in temporal difference error, which further effectively manages risk exposure in potential profit losses. Case studies demonstrate superior performance in balancing risk and opportunity compared to benchmarks under varying forecast conditions.
@@ 1. Introduction
Energy price arbitrage have become the main service for grid-scale batteries, with more than 15 GW deployed in California (California Energy Commission, 2025) and 6 GW in Texas (Reuters, 2025). Price arbitrage of energy storage fundamentally relies on accurate predictions of future market conditions and opportunity values. However, due to the intermittency of renewable energy sources, demand fluctuations, network congestions, and complex market dynamics, electricity markets exhibit high volatility, making accurate forecasting extremely challenging. Prediction errors in arbitrage decisions not only reduce immediate profit but also propagate through sequential operations, as current charging or discharging choices affect the future state of charge (SoC) and available actions. The decision value is only revealed retrospectively after the full arbitrage window, making real-time algorithm calibration particularly challenging. These issues significantly impact the economic viability of energy storage and underscore the need for uncertainty-aware, sequential decision-making frameworks.
Existing approaches to managing risk and prediction uncertainty in storage arbitrage remain inadequate. Stochastic and robust optimization methods often rely on restrictive distributional assumptions or lead to overly conservative strategies that forgo profitable opportunities (Wu et al., 2025). Online learning approaches require real-time feedback, which is unavailable in arbitrage settings. In contrast, conformal prediction and conformal decision theory provide distribution-free uncertainty quantification with formal coverage guarantees (Gibbs and Candes, 2021; Angelopoulos and Bates, 2022; Angelopoulos et al., 2023a; Feldman et al., 2023; Angelopoulos et al., 2023b; Lekeufack et al., 2024; Li and Rodríguez, 2025), enabling decision-makers to explicitly account for prediction uncertainty without assuming specific probability distributions. While recent studies have explored conformal calibration and risk-aware training frameworks (Yeh et al., 2024, 2025), they focus on static, day-ahead settings (Donti et al., 2017) and overlook the temporal dependencies and evolving uncertainty of sequential arbitrage. Further research is needed to extend conformal methods for real-time, adaptive energy storage arbitrage under uncertainty.
This paper addresses the critical challenge of managing imperfect predictions in online energy storage arbitrage when profit loss feedback is unobservable by developing a novel conformal risk-aware decision-making approach. This work makes the following key contributions:
@@ 2.1. Risk-Aware Energy Storage Arbitrage
Traditional energy storage arbitrage strategies typically adopt a risk-neutral objective, aiming to maximize expected profit using approaches such as stochastic optimization (Krishnamurthy et al., 2017), stochastic dynamic programming (Zheng et al., 2022; Mégel et al., 2015), reinforcement learning (Wang and Zhang, 2018), and hybrid methods that combine machine learning with optimization (Baker et al., 2023; Wu et al., 2024; Yi et al., 2025a, b). These risk-neutral strategies can effectively capture theoretical arbitrage opportunities and achieve near-optimal performance under accurate forecasts. Yet, electricity prices, which do not follow any analytical stochastic process, are inherently volatile due to network congestion and the need to maintain a balance between instantaneous supply and demand. Consequently, electricity price forecasts are often inaccurate. As a result, energy storage arbitrage participants are highly exposed to downside risks arising from imperfect predictions, and incorporating risk considerations into arbitrage optimization is critical.
Recent work in storage arbitrage under uncertainty has explored various risk metrics. These approaches either rely on assumptions about probability distributions of market uncertainties or produce overly conservative strategies. For example, traditional risk management approaches, such as Conditional Value-at-Risk (CVaR) (Jabr, 2005), rely heavily on scenario generation techniques, but are inconvenient for practical implementations given the high volatility and the non-stationary nature of electricity prices. Stochastic optimization approaches, such as robust and chance-constrained optimization methods, while providing worst-case guarantees, tend to be overly conservative and may miss profitable opportunities (Wu et al., 2025). Online uncertainty management approaches, such as online convex optimization (Hazan, 2022; Chen et al., 2015), require observable ground truth or performance feedback, but in energy storage arbitrage, the true value of decisions only becomes apparent retrospectively after the entire arbitrage window, making real-time algorithm adaptation extremely challenging. Moreover, the performance of such algorithms can be further compromised due to storage operational constraints. A summary of these representative algorithms is provided in Table 1.
@@
Risk Neutral $\times$ $\checkmark$ $\times$ (Zheng et al., 2022; Baker et al., 2023)
CVaR $\checkmark$ $\times$ $\times$ (Jabr, 2005)
Chance Constrained $\checkmark$ $\times$ $\times$ (Wu et al., 2025)
Robust Optimization $\checkmark$ $\times$ $\times$ (Wu et al., 2025)
Switching Cost $\checkmark$ $\checkmark$ $\times$ (Chen et al., 2015)
@@ 2.2. Conformal Prediction and Decision Theory
Conformal prediction provides distribution-free uncertainty quantification by constructing prediction sets (a set of possible predictions) with finite-sample coverage guarantees (Papadopoulos et al., 2002; Vovk et al., 2005; Lei and Wasserman, 2013; Angelopoulos and Bates, 2022). A brief summary of conformal prediction is provided in Appendix A for reference. Unlike traditional statistical methods that rely on distributional assumptions, conformal prediction only requires the exchangeability of data points, making it broadly applicable across different domains. Conformal risk control (Angelopoulos et al., 2023a) extends conformal prediction to control arbitrary bounded loss functions beyond simple miscoverage. This framework enables risk control for complex decision-making scenarios where traditional accuracy metrics may not capture the full cost of incorrect decisions.
Another line of work (Gibbs and Candes, 2021; Feldman et al., 2023; Angelopoulos et al., 2023b; Lekeufack et al., 2024) circumvents the assumptions on data exchangeability by adopting a control perspective that dynamically adapts prediction sets to real-time errors to achieve desired target bounds, extending the framework to online settings (Li and Rodríguez, 2025). Recent developments have extended conformal prediction to online settings and sequential decision-making applications. Adaptive Conformal Inference (ACI) (Gibbs and Candes, 2021) enables real-time adaptation to distribution shifts by dynamically adjusting prediction set sizes based on recent coverage performance. Rolling Risk Control (Feldman et al., 2023) extends ACI to control arbitrary bounded loss functions. The conformal PID controller (Angelopoulos et al., 2023b) bridges conformal prediction and control theory, enabling more stable coverage in an online adaptive setting. Recent work on conformal decision theory (Lekeufack et al., 2024) further generalizes these concepts establishing the theoretical foundations for safe autonomous decisions under imperfect predictions.
3. Problem Formulation
@@ 3.1. Storage Arbitrage
We consider the online arbitrage problem of an automated energy storage unit operating in the real-time electricity market. At each time step $t\in\mathcal{T}$, where $\mathcal{T}=\{1,\ldots,T\}$, the unit must decide whether to charge, discharge, or remain idle, with the objective of maximizing the total profit comprised of both the immediate payoff and the expected maximized profit over the remaining arbitrage window. We denote $Q_{t}(e_{t})$ as the opportunity value function of the battery’s SoC $e_{t}\in[0,E]$ at time step $t$. $Q_{t}(e_{t})$ is monotonically non-decreasing in $e_{t}$. Note that the true value of $Q_{t}(e_{t})$ is unknown in real time and only revealed retrospectively after the entire arbitrage window, hence we must resort to predictions for online decision-making. We denote $\hat{Q}_{t}(e_{t}|\lambda_{t})$ as the predicted opportunity value function with respect to SoC $e_{t}$ given the realization of price $\lambda_{t}$ (the detailed prediction method is described in Section 3.2). The optimal arbitrage problem takes the form:
@@
where the first term in the objective function (1a) represents the arbitrage revenue, calculated as the product of the real-time market price $\lambda_{t}$ and the net power dispatch $(p_{t}-b_{t})$, with $p_{t}$ representing discharge power and $b_{t}$ representing charge power. The second term accounts for discharge costs, where $C$ represents the marginal operational cost, including factors such as battery degradation. The third term $\hat{Q}_{t}(e_{t}|\lambda_{t})$ is the predicted opportunity value to go. Constraint (1b) enforces the power rating limit $P$ for both charging and discharging operations. Constraint (1c) provides a relaxed formulation that prevents simultaneous charging and discharging (Xu et al., 2020). Constraint (1d) models the SoC evolution constraint with charging/discharging efficiency $\eta$. Constraint (1e) ensures the SoC remains within the storage capacity limits. To ease notation we define $\mathcal{E}(e_{t-1})$ as the set $(p_{t},b_{t},e_{t})$ such that constraints (1b)-(1e) are feasible given $e_{t-1}$.
@@
By adopting a dynamic programming approach (detailed in Section 3.2), it can be seen that marginal opportunity value functions $q_{t}(e_{t})$ and thus $\hat{q}_{t}(e_{t}|\lambda_{t})$ are non-increasing in $e_{t}$, which makes the objective function (1a) concave in $e_{t}$. It is assumed that the feasible set $\mathcal{E}(e_{t-1})$ is nonempty. Therefore, the optimization problem (1) is convex with a concave-maximizing objective function and non-empty feasible set (Boyd and Vandenberghe, 2004). The predicted marginal opportunity value function, $\hat{q}_{t}(e_{t}|\lambda_{t})$, crucially appears in the analytical solutions to (1), i.e., the storage dispatch decisions $(p^{*}_{t},b^{*}_{t})$ for all $t\in\mathcal{T}$. These solutions can be obtained with the realized price $\lambda_{t}$, the prediction of marginal opportunity value function and current SoC $e_{t-1}$ (Zheng et al., 2022). Detailed formulation and interpretations of the solution are provided in Appendix B. Finally, we denote the dispatch decision at time $t$ as $D_{t}:=(p^{*}_{t},b^{*}_{t})$.
@@ 3.2. Opportunity Value Function Forecasting
The prediction of the marginal opportunity value function $\hat{q}_{t}(e_{t})$ can be obtained through a forecasting model proposed in (Baker et al., 2023), which we briefly review here. The development of this forecasting model involves two key steps: opportunity valuation and model training. First, notice that the optimal arbitrage problem (1) can be formulated within a dynamic programming framework, where the total value given time step represents the maximum achievable arbitrage profit comprising both the payoff from the current time step (the first two terms in (1a)) and the potential opportunity values from future operations (the third term in (1a)). Thus, given a series of prices $\{\lambda_{t}\}_{t\in\mathcal{T}}$, this opportunity value function can be calculated in a backwards recursive manner, as follows:
@@
This approach follows the Bellman equation in dynamic programming (Bellman, 1957) and fits a piece-wise linear approximation of the marginal opportunity value function $q_{t}(e_{t})$ based on the first-order derivative of the optimal value function $Q_{t}(e_{t})$, where $e_{t}$ is drawn from the set of SoC segments $\mathcal{S}$. The valuation of $q_{t}(e_{t})$ can be efficiently computed using analytical solutions in Appendix (17). Once the opportunity value function is calculated, the forecasting model can be trained using historical (and current) prices $\Lambda_{t}$ over a look-back window and the corresponding marginal opportunity value function $q_{t}(e_{t})$. We denote this model as $\mathcal{M}:\Lambda\rightarrow\hat{q}$, where $\Lambda$ is the set $\{\Lambda_{t}\}_{t\in\mathcal{T}}$ and $\hat{q}$ is $\{\hat{q}_{t}(e_{t}|\lambda_{t})\}_{t\in\mathcal{T}}$.
@@ 3.3. Decision Evaluation under Uncertainty
Policy evaluation for online decision-making is non-trivial since decisions are made under uncertainty about future states, making the true quality of current actions unobservable until the outcomes fully unfold. In the context of storage arbitrage, the main objective of policy evaluation is to quantify the (prediction) value error $\Delta(e_{t})~:=~\hat{Q}_{t}(e_{t}|\lambda_{t})~-~Q_{t}(e_{t})$ which represents the distance from the forecast to the true value function. However, $\Delta(e_{t})$ is often unmeasurable as the true value function $Q_{t}(e_{t})$ is unobtainable in real time. To analyze this $\Delta_{t}$, we first reformulate the opportunity valuation (3) into an online framework as follows:
@@
An alternative approximation for value error is the Bellman error (Sutton and Barto, 1998). Under deterministic environment and policy conditions, as in our setup with transition dynamics and policy defined by (4), this Bellman error is identical to the temporal difference error (Sutton, 1988), defined as:
@@
The Bellman equation has a unique solution given by the true value function (Bertsekas and Tsitsiklis, 1996). Alternatively, the solution of the Bellman equation can be viewed from a temporal-difference perspective in that it is analogous to the temporal difference error being zero. Finally, let $\Pi_{t}$ denote the step arbitrage profit as the sum of first two terms in (1a).
Proposition 0.
Proof.
$\Delta(e_{t})$ $=\hat{Q}_{t}(e_{t}|\lambda_{t})-Q_{t}(e_{t})$
$=\hat{Q}_{t}(e_{t}|\lambda_{t})-(\Pi_{t}+Q_{t}(e_{t+1}))$
$(\Rightarrow)$ First we show that if $\delta(e_{t})=0$, then $\Delta(e_{t})=0$, for all $e_{t}\in[0,E],\ t\in\mathcal{T}$. By assumption, we have the boundary condition, at $t=T$, $\Delta(e_{T})=0$ for all $e_{T}\in[0,E]$. Assume $\Delta(e_{t+1})=0$ for all $e_{t+1}\in[0,E]$, then for time step $t$, using the recursive relationship, $\Delta(e_{t})=\delta(e_{t})+\Delta(e_{t+1})=0$ since $\delta(e_{t})=0$. Therefore, by induction, $\Delta(e_{t})=0$ if $\delta(e_{t})=0$ for all $e_{t}\in[0,E],\ t\in\mathcal{T}$.
Proposition 1 suggests that the temporal difference error serves as a measurable proxy for value error, which is the true but typically unobservable objective. Due to the existence of such errors, sequential online decisions under opportunity value function forecasting inevitably yield suboptimal results with respect to perfect information. In the context of storage arbitrage, this value error manifests as a profit loss relative to the theoretical optimum, as further discussed in Remark 2.
@@
Remark 2 (Inevitable profit loss under imperfect predictions).
Following Bellman’s principle of optimality (Bellman, 1957), the solution to the arbitrage problem (3) yields the optimal storage dispatch decisions that maximize the total profit over the arbitrage window. That is, when opportunity value functions are estimated accurately, i.e., $\hat{Q}_{t}(e_{t})={Q}_{t}(e_{t})$ for all $e_{t}\in[0,E],\ t\in\mathcal{T}$, the calculated per-step decisions are globally optimal. In this ideal case, it’s easy to show that the total value—incurred profits plus remaining opportunity value—remains constant throughout the arbitrage window. A strategy achieving this is said to fully capture and realize the theoretical opportunity value. The ultimate objective of a risk-aware arbitrage strategy is to constrain the gap between the theoretical optimum and realized discounted profits, i.e., profit loss in hindsight. Such a gap is often referred to as regret in the online algorithm literature. Given a certain sequence of decision decisions (dispatch policy) $\mathcal{D}_{1:T}:=\{(p_{t},b_{t})\}_{t\in\mathcal{T}}$, we formally define this as follows:

### v2

@@ Abstract.
This work proposes a conformal approach for energy storage arbitrage to control the downside risk arising from imperfect price forecasts. Energy storage arbitrage relies solely on predictions of future market prices, while inaccurate price predictions may lead to significant profit losses. Based on conformal decision theory, we describe a controller that dynamically adjusts decision conservativeness through prediction sets without distributional assumptions. To enable online calibration when online profit loss feedback is unobservable, we establish that a temporal difference error serves as a measurable proxy. Building on this insight, we develop two online calibration strategies: prediction error-based adaptation targeting forecast accuracy, and value error-based calibration focusing on decision quality. Analysis of the conformal controller proves bounded long-term risk with convergence guarantees in temporal difference error, which further effectively manages risk exposure in potential profit losses. Case studies demonstrate superior performance in balancing risk and opportunity compared to benchmarks under varying forecast conditions.
@@ 1. Introduction
Energy price arbitrage has become the main service for grid-scale batteries, with more than 15 GW deployed in California (CaliforniaStorage2024) and 6 GW in Texas (TexasStorage2025). Price arbitrage of energy storage fundamentally relies on accurate predictions of future market conditions and opportunity values. However, due to the intermittency of renewable energy, demand fluctuations, network congestion, and complex market dynamics, electricity markets exhibit high volatility, making accurate forecasting extremely challenging. Prediction errors in arbitrage decisions not only reduce immediate profit but also propagate through sequential operations, as current charging or discharging choices affect the future state of charge (SoC) and available actions. The decision value is only revealed retrospectively after the full arbitrage window, for instance, when evaluating annual profit once all prices within the year are realized, making real-time algorithm calibration particularly challenging. These issues significantly impact the economic viability of energy storage and underscore the need for uncertainty-aware, sequential decision-making frameworks.
Existing approaches to managing risk and prediction uncertainty in storage arbitrage remain inadequate. Stochastic and robust optimization methods often rely on restrictive distributional assumptions or lead to overly conservative strategies that forgo profitable opportunities (wu2025). Online learning approaches require real-time feedback, which is unavailable in arbitrage settings. In contrast, conformal prediction and conformal decision theory provide distribution-free uncertainty quantification with formal coverage guarantees (gibbs2021adaptive; angelopoulos2022; angelopoulos2023risk; feldman2023achieving; angelopoulos2023; lekeufack2024; li2025neural), enabling decision-makers to explicitly account for prediction uncertainty without assuming specific probability distributions. While recent studies have explored conformal calibration and risk-aware training frameworks (yeh2024; yeh2025), they focus on static, day-ahead settings (donti2017) and overlook the temporal dependencies and evolving uncertainty of sequential arbitrage. Further research is needed to extend conformal methods for real-time, adaptive energy storage arbitrage under uncertainty.
This paper addresses the critical challenge of managing imperfect predictions in online energy storage arbitrage when profit loss feedback is unobservable by developing a novel conformal risk-aware decision-making approach. This work makes the following contributions:
@@ 2.1. Risk-Aware Energy Storage Arbitrage
Traditional energy storage arbitrage strategies typically adopt a risk-neutral objective, aiming to maximize expected profit using approaches such as stochastic optimization (krishnamurthy2017energy), stochastic dynamic programming (zheng2022; megel2015stochastic), reinforcement learning (wang2018energy), and hybrid methods that combine machine learning with optimization (baker2023transferable; wu2024; yi2025; yi2025decision). These risk-neutral strategies can effectively capture theoretical arbitrage opportunities and achieve near-optimal performance under accurate forecasts. Yet, electricity prices, which do not follow any analytical stochastic process, are inherently volatile due to network congestion and the need to maintain a balance between instantaneous supply and demand. Consequently, electricity price forecasts are often inaccurate. As a result, energy storage arbitrage participants are exposed to downside risk arising from imperfect predictions, and incorporating risk considerations into arbitrage optimization is critical.
Recent work in storage arbitrage under uncertainty has explored various risk metrics. These approaches either rely on assumptions about probability distributions of market uncertainties or produce overly conservative strategies. For example, traditional risk management approaches, such as Conditional Value-at-Risk (CVaR) (jabr2005robust), rely heavily on scenario generation techniques, but are inconvenient for practical implementations given the high volatility and the non-stationary nature of electricity prices. Stochastic optimization approaches, such as robust and chance-constrained optimization methods, while providing worst-case guarantees, tend to be overly conservative and may miss profitable opportunities (wu2025). Online uncertainty management approaches, such as online convex optimization (hazan2022introduction; chen2015), require observable ground truth or performance feedback, but in energy storage arbitrage, the true value of decisions only becomes apparent retrospectively after the entire arbitrage window, making real-time algorithm adaptation extremely challenging. Moreover, the performance of such algorithms can be further compromised due to storage operational constraints. A summary of these representative algorithms is provided in Table 1.
@@
Risk Neutral $\times$ $\checkmark$ $\times$ (zheng2022; baker2023transferable)
CVaR $\checkmark$ $\times$ $\times$ (jabr2005robust)
Chance Constrained $\checkmark$ $\times$ $\times$ (wu2025)
Robust Optimization $\checkmark$ $\checkmark$ $\times$ (wu2025)
Switching Cost $\checkmark$ $\checkmark$ $\times$ (chen2015)
@@ 2.2. Conformal Prediction and Decision Theory
Conformal prediction provides distribution-free uncertainty quantification by constructing prediction sets (a set of possible predictions) with finite-sample coverage guarantees (papadopoulos2002; vovk2005; lei2013; angelopoulos2022). A brief summary of conformal prediction is provided in Appendix A for reference. Unlike traditional statistical methods that rely on distributional assumptions, conformal prediction only requires the exchangeability of data points, making it broadly applicable across different domains. Conformal risk control (angelopoulos2023risk) extends conformal prediction to control arbitrary bounded loss functions beyond simple miscoverage. This framework enables risk control for complex decision-making scenarios where traditional accuracy metrics may not capture the full cost of incorrect decisions.
Another line of work (gibbs2021adaptive; feldman2023achieving; angelopoulos2023; lekeufack2024) circumvents the assumptions on data exchangeability by adopting a control perspective that dynamically adapts prediction sets to real-time errors to achieve desired target bounds, extending the framework to online settings (li2025neural). Recent developments have extended conformal prediction to online settings and sequential decision-making applications. Adaptive Conformal Inference (ACI) (gibbs2021adaptive) enables real-time adaptation to distribution shifts by dynamically adjusting prediction set sizes based on recent coverage performance. Rolling Risk Control (feldman2023achieving) extends ACI to control arbitrary bounded loss functions. The conformal PID controller (angelopoulos2023) bridges conformal prediction and control theory, enabling more stable coverage in an online adaptive setting. Recent work on conformal decision theory (lekeufack2024) further generalizes these concepts establishing the theoretical foundations for safe autonomous decisions under imperfect predictions.
3. Problem Formulation and Preview of Main Results
@@ 3.1. Storage Arbitrage
We consider the online arbitrage problem of an automated energy storage unit participating in the real-time electricity market as a price-taking participant. At each time step $t\in\mathcal{T}$, where $\mathcal{T}=\{1,\ldots,T\}$ defines the arbitrage window, the unit must decide whether to charge, discharge, or remain idle, with the objective of maximizing the total profit comprised of both the immediate payoff and the expected maximized profit over the remaining arbitrage window $\{t+1,\ldots,T\}$. We denote $Q_{t}(e_{t})$ as the opportunity value function of the battery’s SoC $e_{t}\in[0,E]$ at time step $t$. $Q_{t}(e_{t})$ is monotonically non-decreasing in $e_{t}$. Note that the true value of $Q_{t}(e_{t})$ is unknown in real time and only revealed retrospectively after the entire arbitrage window, hence we must resort to predictions for online decision-making. We denote $\hat{Q}_{t}(e_{t}|\lambda_{t})$ as the predicted opportunity value function with respect to SoC $e_{t}$ given the realization of price $\lambda_{t}$ (the detailed prediction method is described in Section 3.2). The optimal arbitrage problem takes the form:
@@
where the first term in the objective function (1a) represents the arbitrage revenue, calculated as the product of the real-time market price $\lambda_{t}$ and the net power dispatch $(p_{t}-b_{t})$, with $p_{t}$ representing discharge power and $b_{t}$ representing charge power. The second term accounts for discharge costs, where $C$ represents the marginal operational cost, including factors such as battery degradation. The third term $\hat{Q}_{t}(e_{t}|\lambda_{t})$ is the predicted opportunity value to go. Constraint (1b) enforces the power rating limit $P$ for both charging and discharging operations. Constraint (1c) provides a relaxed formulation that prevents simultaneous charging and discharging (xu2020). This assumption is reasonable since discharging at negative prices is economically irrational. Constraint (1d) models the SoC evolution constraint with charging/discharging efficiency $\eta$. Constraint (1e) ensures the SoC remains within the storage capacity limits. To ease notation we define $\mathcal{E}(e_{t-1})$ as the set $(p_{t},b_{t},e_{t})$ such that constraints (1b)-(1e) are feasible given $e_{t-1}$.
@@
By adopting a dynamic programming approach (detailed in Section 3.2), it can be seen that marginal opportunity value functions $q_{t}(e_{t})$ and thus $\hat{q}_{t}(e_{t}|\lambda_{t})$ are non-increasing in $e_{t}$, which makes the objective function (1a) concave in $e_{t}$. It is assumed that the feasible set $\mathcal{E}(e_{t-1})$ is nonempty. Therefore, the optimization problem (1) is convex with a concave-maximizing objective function and non-empty feasible set (boyd2004). The predicted marginal opportunity value function, $\hat{q}_{t}(e_{t}|\lambda_{t})$, crucially appears in the analytical solutions to (1), i.e., the storage dispatch decisions $(p^{*}_{t},b^{*}_{t})$ for all $t\in\mathcal{T}$. These solutions can be obtained with the realized price $\lambda_{t}$, the prediction of marginal opportunity value function and current SoC $e_{t-1}$ (zheng2022). Detailed formulation and interpretations of the solution are provided in Appendix B. Finally, we denote the dispatch decision at time $t$ as $D_{t}:=(p^{*}_{t},b^{*}_{t})$.
Remark 1.
Storage arbitrage problem (1) applies to both price-taker market bidding and price response modes. The key distinction lies in the time of decision-making relative to market price realization.
  1. (1)
    Market bidding: storage submits its bids a predetermined time before the market clearing period $t$ and follows the dispatch results determined by market clearing. Given that bids for time step $t$ are designed indeed based on the price forecast for $\{t+1,\ldots,T\}$ and market clearing relies only on current bids, the joint problem of optimal bid design and market clearing is equivalent to the formulation (1(baker2023transferable; yi2025decision).
  2. (2)
    Price response: storage updates its dispatch decisions at each time step $t$ after observing the most recent real-time market price $\lambda_{t}$, announced by the system operator prior to dispatch.
@@ 3.2. Opportunity Value Function Forecasting
The prediction of the marginal opportunity value function $\hat{q}_{t}(e_{t})$ can be obtained through a forecasting model proposed in (baker2023transferable), which we briefly review here. The development of this forecasting model involves two key steps: opportunity valuation and model training. First, notice that the optimal arbitrage problem (1) can be formulated within a dynamic programming framework, where the total value given time step represents the maximum achievable arbitrage profit comprising both the payoff from the current time step (the first two terms in (1a)) and the potential opportunity values from future operations (the third term in (1a)). Thus, given a series of prices $\{\lambda_{t}\}_{t\in\mathcal{T}}$, this opportunity value function can be calculated in a backwards recursive manner, as follows:
@@
This approach follows the Bellman equation in dynamic programming (bellman1957dynamic) and fits a piece-wise linear approximation of the marginal opportunity value function $q_{t}(e_{t})$ based on the first-order derivative of the optimal value function $Q_{t}(e_{t})$, where $e_{t}$ is drawn from the set of SoC segments $\mathcal{S}$. The valuation of $q_{t}(e_{t})$ can be efficiently computed using analytical solutions in Appendix (18). Once the opportunity value function is calculated, the forecasting model can be trained using historical (and current) day-ahead and real-time prices $\Lambda_{t}$ over a look-back window and the corresponding marginal opportunity value function $q_{t}(e_{t})$. We denote this model as $\mathcal{M}:\Lambda\rightarrow\hat{q}$, where $\Lambda$ is the set $\{\Lambda_{t}\}_{t\in\mathcal{T}}$ and $\hat{q}$ is $\{\hat{q}_{t}(e_{t}|\lambda_{t})\}_{t\in\mathcal{T}}$.
@@ 3.3. Decision Evaluation under Uncertainty
Policy evaluation for online decision-making is non-trivial since decisions are made under uncertainty about future states, making the true quality of current actions unobservable until the outcomes fully unfold. In the context of storage arbitrage, the main objective of policy evaluation is to quantify the (prediction) value error:
(4) $\Delta(e_{t}):=\hat{Q}_{t}(e_{t}|\lambda_{t})-Q_{t}(e_{t})$
which represents the distance from the forecast to the true value function. However, $\Delta(e_{t})$ is often unmeasurable as the true value function $Q_{t}(e_{t})$ is unobtainable in real time. To analyze this $\Delta_{t}$, we first reformulate the opportunity valuation (3) into an online framework as follows:
@@
An alternative approximation for value error is the Bellman error (sutton1998reinforcement). Under deterministic environment and policy conditions, as in our setup with transition dynamics and policy defined by (5), this Bellman error is identical to the temporal difference error (sutton1988), defined as:
@@
The Bellman equation has a unique solution given by the true value function (bertsekas1996). Alternatively, the solution of the Bellman equation can be viewed from a temporal-difference perspective in that it is analogous to the temporal difference error being zero (fujimoto2022), which will later be formalized in Proposition 3. Due to the existence of such errors, sequential online decisions under opportunity value function forecasting inevitably yield suboptimal results with respect to perfect information. In the context of storage arbitrage, this value error manifests as a profit loss relative to the theoretical optimum, as further discussed in Section 3.4.
@@
3.4. Problem Statement
Following Bellman’s principle of optimality (bellman1957dynamic), the solution to the arbitrage problem (3) yields the optimal storage dispatch decisions that maximize the total profit over the arbitrage window. That is, when opportunity value functions are estimated accurately, i.e., $\hat{Q}_{t}(e_{t})={Q}_{t}(e_{t})$ for all $e_{t}\in[0,E],\ t\in\mathcal{T}$, the calculated per-step decisions are globally optimal. In this ideal case, it’s easy to show that the total value, i.e., incurred profits plus remaining opportunity value, remains constant throughout the arbitrage window. A strategy achieving this is said to fully capture and realize the theoretical opportunity value. The ultimate objective of a risk-aware arbitrage strategy is to constrain the gap between the theoretical optimum and realized discounted profits, i.e., profit loss in hindsight. Such a gap is often referred to as regret in the online algorithm literature. Let $\Pi_{t}$ denote the step arbitrage profit as the sum of first two terms in (1a). Given a certain sequence of decision decisions (dispatch policy) $\mathcal{D}_{1:T}:=\{(p_{t},b_{t})\}_{t\in\mathcal{T}}$, we formally define this as follows: